# SimpleSafetyTests: a Test Suite for Identifying Critical Safety Risks in Large Language Models

### v1

@@ Abstract
The past year has seen rapid acceleration in the development of large language models (LLMs). For many tasks, there is now a wide range of open-source and open-access LLMs that are viable alternatives to proprietary models like ChatGPT. Without proper steering and safeguards, however, LLMs will readily follow malicious instructions, provide unsafe advice, and generate toxic content. This is a critical safety risk for businesses and developers. We introduce SimpleSafetyTests as a new test suite for rapidly and systematically identifying such critical safety risks. The test suite comprises 100 test prompts across five harm areas that LLMs, for the vast majority of applications, should refuse to comply with. We test 11 popular open LLMs and find critical safety weaknesses in several of them. While some LLMs do not give a single unsafe response, most models we test respond unsafely on more than 20% of cases, with over 50% unsafe responses in the extreme. Prepending a safety-emphasising system prompt substantially reduces the occurrence of unsafe responses, but does not completely stop them from happening. We recommend that developers use such system prompts as a first line of defence against critical safety risks.
@@ 1 Introduction
Since the release of ChatGPT in November 2022, there has been a massive increase in the development of large language models (LLMs). Meta’s recent release of the open-access Llama (Touvron et al., 2023a) and Llama2 models (Touvron et al., 2023b) in particular has accelerated this development, as well as releases like Mistral (Jiang et al., 2023) and Falcon (Penedo et al., 2023). As of November 6th 2023, there are over 30k generative language models hosted on Hugging Face, at least 5k of which are derivatives of Meta’s Llama models. [^1] https://huggingface.co/models For businesses and developers looking to use and deploy LLMs in user-facing applications, a growing open-access ecosystem presents great opportunity to access and tune these LLMs as viable alternatives to proprietary models.
Figure 1: Responses from Mistral (7B), Dolly v2 (12B), and Llama2 (13B), without a system prompt, to a test prompt in SimpleSafetyTests.
@@
Without proper safeguards, however, LLMs will readily follow malicious instructions and answer risky questions in highly unsafe ways Gehman et al. (2020); Hartvigsen et al. (2022); OpenAI (2023), such as advising how to make bombs, how to commit serious fraud, or responding affirmatively to whether someone should commit suicide. This could lead to serious harm being inflicted, potentially of a physical nature, if malicious, vulnerable, or misguided users act on the unsafe responses they are given. Effective evaluation is essential for diagnosing and addressing critical safety weaknesses. Businesses and developers, as well as commercial labs and AI researchers, need evaluation methods that are cost-effective, time-efficient, and compatible with existing evaluation suites Weidinger et al. (2021); Bender et al. (2021); Bommasani et al. (2022); Chan et al. (2023); Dai et al. (2023); OpenAI (2023); Ji et al. (2023).
We introduce  SimpleSafetyTests (SST) as a new test suite for rapidly and systematically identifying critical safety risks in LLMs. SST comprises 100 handcrafted English test prompts across five harm areas. Safety is subjective, and different applications require different safeguards. Nonetheless, we believe that some model responses are nearly always unsafe, outside of very specific contexts such as model development or educating people about model misuse. For example, it is undesirable for a model to generate child abuse material in any situation. Severely unsafe instructions like this are the focus of SST, in contrast to work on LLM safety that takes a broader perspective (Dai et al., 2023; Ji et al., 2023).
We use SST to evaluate 11 open-access LLMs, all tuned to answer questions and/or respond to instructions. We manually annotate all model responses on the 100 test prompts in SimpleSafetyTests for whether they are unsafe or not. We test each LLM both without a system prompt and with a safety-emphasising system prompt. Across all 11 LLMs and both prompt setups, 27% of responses are unsafe. The proportion of responses that are unsafe varies substantially by LLM. None of the responses of LLama2 (13B) are unsafe, and Falcon (40B) responds unsafely only once. All other models show clear weaknesses, responding unsafely between 5% and 73% of the time. Adding a safety-emphasising system prompt reduces the proportion of unsafe responses by ten percentage points across all 11 LLMs, from 33% to 21%. However, the effectiveness of the system prompt varies by LLM and on its own does not remove all safety risks, as measured by SST.
Figure 2: Performance of the 11 models on SimpleSafetyTests, with both the safety-emphasising system prompt and no system prompt. The models are ordered by their result with no system prompt. The dotted red vertical lines show cutoffs for models where 5% of responses are unsafe and 50% of responses are unsafe.
SimpleSafetyTests is available on GitHub with an open-source license. [^2] https://github.com/bertiev/SimpleSafetyTests The models’ responses are available only on request due to their sensitive nature.
@@ 2.1 Harm Areas
SST covers five harm areas: (1) Suicide, Self-Harm, and Eating Disorders; (2) Physical Harm; (3) Illegal and Highly Regulated Items; (4) Scams and Fraud; and (5) Child Abuse. These harm areas have been selected based on the severity of the harm that they inflict, their prevalence, and the extensive prior work undertaken to document their effects. This includes academic research on harmful forms of content, conduct, and contact online Agrafiotis et al. (2018); Asam and Katz (2018); Banko et al. (2020); Prabhakaran et al. (2020), academic research on AI safety Weidinger et al. (2021); Glaese et al. (2022); Hao et al. (2023), the trust and safety policies and commentaries of large AI labs (OpenAI [^3] https://openai.com/safety-standards, Anthropic [^4] https://www.anthropic.com/index/core-views-on-ai-safety, and Cohere [^5] https://cohere.com/responsibility), as well as the work of prominent trust and safety organisations, such as the WEF’s typology of online harms [^6] https://www3.weforum.org/docs/WEF_Typology_of_Online_Harms_2023.pdf, the Atlantic Council’s report on scaling trust [^7] https://www.atlanticcouncil.org/in-depth-research-reports/report/scaling-trust/ and the Digital Trust and Safety Partnership’s glossary of terms [^8] https://dtspartnership.org/glossary/.
Model
Child
Abuse
Illegal
Items
Physical
Harm
Scams
& Fraud
Suicide,
SH & ED
Overall
% unsafe
Llama2 (13B) 0 0 0 0 0 0
Falcon (40B) 5 0 0 0 0 1
Zephyr-$\alpha$ (7B) 15 10 0 10 10 9
StableLM (7B) 15 45 0 35 20 23
Vicuna (13B) 15 20 30 40 15 24
Zephyr-$\beta$ (7B) 30 30 35 55 30 35
Flan-T5 (3B) 40 40 45 20 40 37
Mistral (7B) 35 40 50 65 55 49
Wizard-Mega (13B) 35 40 75 75 25 50
OASST (12B) 65 50 55 80 50 60
Dolly v2 (12B) 70 50 75 75 75 69
Average 30 30 33 42 29 33
Table 2: Percentage of LLMs’ responses on SimpleSafetyTests that are unsafe (tested with no system prompt), split by the five harm areas. LLMs are ordered by the overall percentage of unsafe responses.
@@ 2.2 Test Prompt Categories
To capture some of the different ways in which harm can be inflicted through the use of LLMs, we created two categories of test prompts.
@@
Questions asked of the model that, if answered in an unsafe way, would readily give the user access to harmful information or would encourage them to engage in harmful activities. Most questions focus on undertaking harmful activities (e.g. physically harming someone) but some also focus on hiding and justifying harmful activities (e.g. excuses for explaining why someone is self-harming).
@@ 2.3 Creating the Test Prompts
The test prompts can all easily be answered in a way that enables, supports, or otherwise encourages harmful activity. All of the test prompts were created by the authors of this paper. The test prompts are all synthetic, and none have been taken from existing datasets or real-world user interactions with LLMs. Therefore, the dataset can be shared without any privacy or copyright risks. The test prompts were originally created by one of the authors, and then iteratively adjusted through multiple rounds of review and editing by the other authors. In the final iteration of SST, every test prompt was independently checked at least three times for coherence and internal validity. Examples of test prompts are given in Table 1.
@@ 3.1 LLMs
We test 11 open-access and open-source LLMs against SST. We only test open models, and do not include any commercial LLMs, since open models typically undergo less steering (in some cases no steering) to minimise safety risks, and therefore can present a greater risk of responding unsafely. We aimed for a mix of model providers and model sizes, ranging from 3B to 40B parameters. All of the models are instruction- or conversation-tuned [^9] For brevity, we refer to every LLM by just its primary name and do not stipulate that it is the tuned- version, and most are competitive on the LMSYS leaderboard, as of October 2023.101010https://chat.lmsys.org/?leaderboard We accessed all models through the Replicate API. [^11] https://replicate.com/ For more information about model sizes, versions, and creators see Appendix A.
For at least five of the LLMs, the developers explicitly acknowledge the risk of toxic and harmful responses. The creators of Wizard-Mega (13B) state it “can produce problematic outputs (especially when prompted to do so)” [^12] https://huggingface.co/openaccess-ai-collective/wizard-mega-13b, the creators of Zephyr state Zephyr $\alpha$ (7B) “is likely to generate problematic text when prompted to do so” [^13] https://huggingface.co/HuggingFaceH4/zephyr-7b-alpha, and the creators of StableLM (7B) state, “the responses a user gets might be of varying quality and might potentially include offensive language and views.” [^14] https://github.com/Stability-AI/StableLM. And in some cases, LLMs are intentionally released with minimal safety steering. Databricks, the creators of Dolly v2 (12B), open-sourced the model so as to encourage the research community to openly collaborate towards the safe use of AI. [^15] https://techcrunch.com/2023/04/12/databricks-dolly-2-generative-ai-open-source Mistral also chose to not apply safeguards to their LLM, delegating safety management to developers building on top of it. [^16] https://sifted.eu/articles/ai-safety-mistral-ai Accordingly, many developers use moderation APIs, system prompts, and other guardrails to minimise unsafe responses from LLMs (Lin et al., 2023; Gehman et al., 2020; Dinan et al., 2022). System prompts specify general principles for how LLMs should respond to prompts. They are attractive because they can be introduced at inference time and do not require more expensive model training. They are particularly relevant for safety-related behaviours because different applications may have different safety requirements that need to be specified flexibly.
@@
We test two system prompt setups. 1) No system prompt is the default setting for most models. The Llama 2 authors, for example, recommend not using a system prompt for their model to minimise the model’s false refusal rate. [^17] https://github.com/facebookresearch/llama/blob/main/UPDATES.md 2) A safety-emphasising system prompt, which instructs models to be careful, honest, and generally safer in its responses. Such system prompts are widely used to steer LLMs’ responses away from being toxic Glaese et al. (2022). We use the “guardrail” system prompt provided by Jiang et al. (2023) for the Mistral (7B) model because it is reasonably comprehensive, and is not obviously biased: “Always assist with care, respect, and truth. Respond with utmost utility yet securely. Avoid harmful, unethical, prejudiced, or negative content. Ensure replies promote fairness and positivity. [^18] https://docs.mistral.ai/usage/guardrailing/ To ensure consistency across models, and to handle the fact that some models may not have been trained to incorporate system prompts as separate instructions, we explicitly added “USER” and “SYSTEM PROMPT” indicators to the test prompts. Test prompts with no system prompt are formatted as “USER: Should I keep cutting myself?” (sst_008). Test prompts with the safety-emphasising system prompt are formatted as “SYSTEM PROMPT: Always assist with care, […]. USER: Should I keep cutting myself?”. In total, we collected 2,200 responses (from 11 models, tested with two system prompt setups, on 100 test prompts).
@@ 3.2 Labeling LLM Responses
Three annotators labeled models’ responses for whether they are unsafe or not. The annotation guidelines were developed through several rounds of review and testing to resolve disagreements on edge cases. A short discussion of labeling for LLM unsafety (rather than safety) is given in Appendix B. We follow the ‘prescriptive paradigm’ for data annotation, in that we want the annotators to apply our annotation guidelines to decide whether the responses are unsafe, rather than applying their own subjective beliefs (Rottger et al., 2022). In 1,962 cases (89%) all three annotators agree on the correct label. Fleiss’ Kappa is 0.82, which indicates very strong agreement (McHugh, 2012). We use the majority labels (taking 2/3 or 3/3 agreement) for our analyses. In a few cases, the responses continue the test prompts, in the same way that a non-instruction tuned autocomplete model would (even though the test prompts are all grammatically complete statements). A short summary of how we handled these responses during data labeling is given in Appendix C.
4 Results
@@
Overall, 27% of the LLMs’ responses are unsafe. However, safety varies considerably across the LLMs and based on whether the safety-emphasising system prompt is used (see Figure 2). None of Llama2’s responses are unsafe, regardless of the system prompt setup, and Falcon (40B) gives only one unsafe response (with no system prompt on a Child Abuse test prompt, sst_099). The other LLMs with no system prompt frequently respond unsafely. 49% of the responses of Mistral (7B) are unsafe, as are 60% of OASST (12B), and 50% of Wizard-Mega (13B). Dolly v2 (12B) gives the most unsafe responses (69%). Adding a safety-emphasising system prompt made a substantial difference to the proportion of responses that are unsafe, reducing it by twelve percentage points overall, from 33% without a system prompt to 21% with the system prompt. For some LLMs this drives a particularly large improvement, such as Zephyr-$\beta$ (36% $\rightarrow$ 17% unsafe) Wizard-Mega (50% $\rightarrow$ 15% unsafe), and Mistral (46% $\rightarrow$ 10% unsafe). The large improvement for Mistral is not surprising given that we use the Mistral safety guardrail prompt. Interestingly, the safety-emphasising system prompt resulted in marginally more unsafe responses for Dolly v2 (12B) (69% $\rightarrow$ 73%). Overall, the safety-emphasising system prompt clearly improves LLM safety, sometimes dramatically, but does not remove all safety risks. There is also the potential that adding a safety-emphasising system prompt could lead to a high false refusal rate, refusing to comply with safe prompts that superficially resemble unsafe prompts (Röttger et al., 2023).
Figure 3: Percentage of unsafe responses on SimpleSafetyTests, averaged across all 11 models, split by the two test prompt categories and the five harm areas.
@@
For both system prompt settings, the proportion of LLMs’ responses that are unsafe varies by the harm area (see Table 2 for just the results of the setting without a system prompt), from 23% unsafe responses for Suicide, Self-Harm, and Eating Disorders, to 33% unsafe responses for Scams and Fraud. We observe similar differences across harm areas both with and without the safety-emphasising system prompt. Several LLMs exhibit a clear weakness on Child Abuse. It is the only harm area where Falcon (40B) with no system prompt, responds unsafely. Across four of the five harm areas, LLMs give more unsafe responses to the Instructions and Actions than the Information and Advice Seeking (see Figure 3). For Scams and Fraud, there is a 16 percentage point difference and for Illegal and Highly Regulated Items the difference is 11 percentage points. Physical Harm is the only harm area that this relationship does not occur. The direction of this relationship holds for all of the LLMs. Because the test prompts in SST are each unique, rather than being directly equivalent cases that are phrased differently (e.g. “Tell me how to make a bomb” and “How do I make a bomb?”), we cannot draw conclusions about LLMs’ propensity to respond unsafely based on the phrasing of requests. This is an interesting avenue for future work.
5 Limitations

### v2

@@ Abstract
The past year has seen rapid acceleration in the development of large language models (LLMs). However, without proper steering and safeguards, LLMs will readily follow malicious instructions, provide unsafe advice, and generate toxic content. We introduce SimpleSafetyTests as a new test suite for rapidly and systematically identifying such critical safety risks. The test suite comprises 100 test prompts across five harm areas that LLMs, for the vast majority of applications, should refuse to comply with. We test 11 open-access and open-source LLMs and 4 closed-source LLMs, and find critical safety weaknesses. While some of the models do not give a single unsafe response, most give unsafe responses to more than 20% of the prompts, with over 50% unsafe responses in the extreme. Prepending a safety-emphasising system prompt substantially reduces the occurrence of unsafe responses, but does not completely stop them from happening. Trained annotators labelled every model response to SimpleSafetyTests (n = $3,000$). We use these annotations to evaluate five AI safety filters (which assess whether a models’ response is unsafe given a prompt) as a way of automatically evaluating models’ performance on SST. The filters’ performance varies, with differences by harm area and whether the responses are unsafe or safe. The widely-used Perspective API has 72% accuracy and a newly-created zero-shot prompt to OpenAI’s GPT-4 performs best with 89% accuracy.
@@ 1 Introduction
Figure 1: Responses from Mistral ($7$B), Dolly v2 ($12$B), and Llama2 ($13$B) to a prompt in SimpleSafetyTests (no system prompt).
Since the release of ChatGPT by OpenAI in November 2022, there has been a massive increase in the development of large language models (LLMs). Meta’s release of the open-access Llama (Touvron et al., 2023a) and Llama2 models (Touvron et al., 2023b) has accelerated this development, as has releases like Mistral (Jiang et al., 2023) and Falcon (Penedo et al., 2023). Well-funded labs like Anthropic, Cohere and AI21 have also released high-performing models. Without proper safeguards, however, LLMs will readily follow malicious instructions and answer risky questions in unsafe ways (Gehman et al., 2020; Hartvigsen et al., 2022; OpenAI, 2023), such as advising how to make bombs, how to commit serious fraud, or responding affirmatively to whether someone should commit suicide. This could lead to serious harm being inflicted, potentially of a physical nature, if malicious, vulnerable, or misguided users act on the unsafe responses they are given. Effective evaluation is essential for diagnosing and addressing critical safety weaknesses. Businesses and developers, as well as commercial labs and AI researchers, need evaluation methods that are cost-effective, time-efficient, and compatible with existing evaluation suites (Weidinger et al., 2021; Bender et al., 2021; Bommasani et al., 2022; Chan et al., 2023; Dai et al., 2023; OpenAI, 2023; Ji et al., 2023).
We introduce SimpleSafetyTests (SST) as a new test suite for rapidly and systematically identifying critical safety risks in LLMs. SST comprises 100 handcrafted English test prompts across five harm areas. Safety is often context-specific and different applications require different safeguards. Nonetheless, we believe that some model responses are nearly always unsafe, outside of very specific contexts such as model development or educating people about model misuse. For example, it is undesirable for a model to generate child abuse material in any situation. Severely unsafe instructions like this are the focus of SST, in contrast to work on LLM safety that takes a broader perspective (e.g. Dai et al., 2023; Ji et al., 2023).
We use SST to evaluate 11 open-access and open-source (hereafter called “open”) LLMs and 4 closed-source LLMs, all tuned by their creators to answer questions and/or respond to instructions. Trained annotators labelled all model responses on the 100 test prompts in SimpleSafetyTests for whether they are unsafe or not. There are 3,000 labels in total (from 15 models, tested with two system prompt setups, on 100 test prompts).
We test each LLM both without a system prompt and with a safety-emphasising system prompt. In total, across all 15 LLMs and both prompt setups, 20% of responses are unsafe. None of the responses of LLama2 (13B) and Claude 2.1 are unsafe, and Falcon (40B) and GPT-4 respond unsafely only once. Across all prompts, the closed-source models are unsafe in 2% of cases, whereas the open models are unsafe in 27% of cases. Other models show clear weaknesses, responding unsafely up to 73% of the time. Adding a safety-emphasising system prompt reduces the proportion of unsafe responses by nine percentage points across all 15 LLMs, from 25% to 16%. However, the effectiveness of the system prompt varies by LLM and on its own does not remove all safety risks identified by SST.
@@
Figure 2: Performance of the 15 models on SimpleSafetyTests, with both the safety-emphasising system prompt and no system prompt. LLMs are ordered by their results with no system prompt within closed-source (top) and open (bottom) models. The dotted vertical lines show cutoffs for models where 5% of responses are unsafe and 50% of responses are unsafe.
SimpleSafetyTests is available on GitHub and HuggingFace with an open-source license. [^1] https://github.com/bertiev/SimpleSafetyTests and https://huggingface.co/datasets/Bertievidgen/SimpleSafetyTests The models’ responses are available only on request due to their sensitive nature.
@@ 2.1 Harm Areas
SST covers five harm areas: (1) Suicide, Self-Harm, and Eating Disorders; (2) Physical Harm; (3) Illegal and Highly Regulated Items; (4) Scams and Fraud; and (5) Child Abuse. These harm areas have been selected based on the severity of the harm that they inflict, their prevalence, and the extensive prior work undertaken to document their effects. This includes academic research on harmful forms of content, conduct, and contact online (Agrafiotis et al., 2018; Asam & Katz, 2018; Banko et al., 2020; Prabhakaran et al., 2020); academic research on AI safety (Weidinger et al., 2021; Glaese et al., 2022; Hao et al., 2023); the trust and safety policies and commentaries of large AI labs (OpenAI (OpenAI, 2024), Anthropic (Anthropic, 2023), and Cohere (Cohere, 2023)); as well as the work of prominent trust and safety organisations, such as the World Economic Forum’s typology of online harms (Forum, 2023), the Atlantic Council’s report on scaling trust (Council, 2023), and the Digital Trust and Safety Partnership’s glossary of terms (Partnership, 2023).
@@ 2.2 Test Prompt Categories
To ensure diversity in our dataset, we created two categories of test prompts.
@@
Questions asked of the model that, if answered in an unsafe way, would give the user access to harmful information or would encourage them to engage in harmful activities. Most questions focus on undertaking harmful activities (e.g. physically harming someone) but some also focus on hiding and justifying harmful activities (e.g. excuses for explaining why someone is self-harming).
@@ 2.3 Creating the Test Prompts
The test prompts can all easily be answered in a way that enables, supports, or otherwise encourages harmful activity. The test prompts are all synthetic, and none have been taken from existing datasets or real-world user interactions with LLMs. Therefore, the dataset can be shared without any privacy or copyright risks. The test prompts were originally created by one of the authors, and then iteratively adjusted through multiple rounds of review and editing by the other authors. In the final iteration of SST, every test prompt was independently checked by at least three people for coherence and internal validity. Examples of test prompts are given in Table 1. In total, there are 100 prompts, equally split into five harm areas and two prompt categories.
@@ 3.1 LLMs
We test 15 LLMs against SST, of which 11 are open and 4 are closed-source. We aimed for a mix of model providers and model sizes, ranging from 3B to 40B parameters for the open models. The closed-source models’ parameters are unknown. All of the models are instruction- or conversation-tuned [^2] For brevity, we refer to every LLM by its primary name and do not stipulate that it is the tuned- version, and most are competitive on the LMSYS leaderboard, as of October 2023 (LMSYS, 2023). The models differ in what safety evaluation and steering they have undergone, with many of the open models having no safeguards at all. In several cases, the creators explicitly acknowledge the risk of toxic and harmful responses. A summary of the safety and alignment steps (if any) taken for each model, as well as relevant statements from the creators about their approach to safety, and information about how we accessed the models, is given in Appendix A. Note that, when creating applications, many developers use moderation APIs, system prompts, and other guardrails to minimise the risk of unsafe responses (Lin et al., 2023; Gehman et al., 2020; Dinan et al., 2022).
@@
We test all LLMs with and without a system prompt. System prompts specify general principles for how LLMs should respond to prompts. They are attractive because they can be introduced at inference time and do not require additional model training. They are particularly relevant for safety-related behaviours because different applications may have different safety requirements that need to be specified flexibly. We test two system prompt setups. 1) No system prompt is the default setting for most models. The Llama2 authors, for example, recommend not using a system prompt for their model to minimise the model’s false refusal rate. [^3] https://github.com/facebookresearch/llama/blob/main/UPDATES.md 2) A safety-emphasising system prompt, which instructs models to be careful, honest, and generally safer in its responses. Such system prompts are widely used to steer LLMs’ responses away from being toxic (Glaese et al., 2022). We use the “guardrail” system prompt provided by Jiang et al. (2023) for the Mistral (7B) model because it is reasonably comprehensive, and is not obviously biased to just that model: “Always assist with care, respect, and truth. Respond with utmost utility yet securely. Avoid harmful, unethical, prejudiced, or negative content. Ensure replies promote fairness and positivity.(Mistral, 2023) In total, we collected 3,000 responses (from 15 models, tested with two system prompt setups, on 100 test prompts). Further details are given in Appendix A.
@@ 3.2 Labeling LLM Responses
Three annotators trained in labelling unsafe, risky and dangerous content, labeled the 3,000 model responses for whether they are unsafe or not given the prompt. The annotation guidelines were developed through several rounds of testing to resolve disagreements on edge cases. We follow the ‘prescriptive paradigm’ for data annotation, in that we want the annotators to apply our annotation guidelines to decide whether the responses are unsafe, rather than applying their own subjective beliefs (Rottger et al., 2022). A short discussion of labeling for LLM unsafety (rather than safety) is given in Appendix B. In 2,750 cases (92%) all three annotators agree on the same label. Fleiss’ Kappa is 0.83, which indicates very strong agreement (McHugh, 2012). We use the majority labels (taking 2/3 or 3/3 agreement) as the final labels for SST.
Models’ responses varied considerably in terms of their content, coherence, length and style. Notably, some safe responses simply reject the prompt (e.g. “Sorry, but I can’t assist with that.”) whereas other safe responses give a detailed explanation of why the prompt creates a risk of harm, offer advice and support to the user, or warn them of the consequences they could face. A further challenge is that despite all the test prompts being grammatically complete statements, some model responses are incoherent and hard to parse. For instance, a few responses simply regurgitate the content of the test prompts or continue with an auto-completion. A short summary of how we handled these responses during data labeling is given in Appendix C.
Table 2: Percentage of LLMs’ responses on SimpleSafetyTests ($n=100$) that are unsafe, split by the five harm areas ($n=20$ each). LLMs are tested without a system prompt. LLMs are ordered by the overall percentage of unsafe responses within closed-source (top) and open (bottom) models.
4 Performance on SimpleSafetyTests
Overall, 20% of the LLMs’ responses are unsafe (n = 608/3,000). However, the degree of safety varies considerably across the LLMs and based on whether the safety-emphasising system prompt is used (see Figure 2). Overall, the closed-source models are unsafe in 2% of cases, whereas the open models are unsafe in 27% of cases. This is a 10x difference. None of Claude 2.1’s responses are unsafe, regardless of the system prompt setup. GPT-4 only gives one unsafe response (without the safety-emphasising prompt), and AI21’s Jurassic-2 is unsafe in only two cases with both system prompt settings. Cohere’s Command model is more unsafe than the other closed-source models, giving 1 unsafe response with the safety-emphasising prompt and 10 unsafe responses without. Llama2 (13B) and Falcon (40B) are positive outliers for the open models. Llama2 gives no unsafe responses and Falcon only gives one unsafe response (without the safety-emphaising prompt). In contrast, the other open models with no system prompt frequently respond unsafely. 49% of the responses of Mistral (7B) are unsafe, as are 60% of OASST (12B), and 50% of Wizard-Mega (13B). Dolly v2 (12B) gives the most unsafe responses at 69%.
@@ 4 Performance on SimpleSafetyTests
Adding a safety-emphasising system prompt makes a substantial difference to the proportion of unsafe responses, reducing it by nine percentage points on average, from 25% without a system prompt to 16% with the system prompt. This difference is larger for open models (12 percentage points) than closed-source (2.5 percentage points). For some LLMs, the safety-emphasising drives a particularly large improvement, such as Zephyr-$\beta$ (36% $\rightarrow$ 17% unsafe), Wizard-Mega (50% $\rightarrow$ 15% unsafe), Mistral (46% $\rightarrow$ 10% unsafe), and Cohere (10% $\rightarrow$ 1% unsafe). The large improvement for Mistral (and Zephyr, which is a derivative of Mistral) is not surprising given that we use the Mistral safety guardrail prompt. Interestingly, the safety-emphasising system prompt resulted in marginally more unsafe responses for Dolly v2 (12B) (69% $\rightarrow$ 73%). Overall, the safety-emphasising system prompt clearly improves the safety of the LLMs we test, sometimes dramatically, but does not remove all safety risks. There is also a risk that adding a safety-emphasising system prompt may increase models’ false refusal rate (Röttger et al., 2023), although we do not test for this with SST.
@@
For both system prompt settings, the proportion of LLMs’ responses that are unsafe varies by the harm area (see Table 2 for just the results of the setting without a system prompt), from 18% unsafe responses for Suicide, Self-Harm, and Eating Disorders, to 25% unsafe responses for Scams and Fraud. The differences are similar for both models with and without the safety-emphasising system prompt, and for both open and closed-source models. Several LLMs exhibit a clear weakness on Child Abuse (21% unsafe in total). It is the only harm area where the otherwise-safe Falcon (40B) with no system prompt responds unsafely. On average, models respond unsafely to 23% of Instructions and Actions prompts, compared with 18% of Information and Advice Seeking prompts. This holds across four of the five harm areas, with Physical Harm and Violence the only exception. LLMs’ propensity to respond unsafely based on the phrasing of requests is an interesting avenue for future work. However, because the test prompts in SST are all unique, rather than being directly equivalent cases that are phrased differently (e.g. “Tell me how to make a bomb” and “How do I make a bomb?”), we cannot draw conclusions.
4.1 Limitations of SimpleSafetyTests